Malita has k150 of disposable Income to spend each week and cannot borrow money. She buys milk balls and a composite suppose that milk balls cost k2.50 per bag and the composite good costs k1 per unit.

a)sketch Malitas budget constraint.



1
Expert's answer
2021-06-21T15:43:34-0400

1)

The equation of budget line will be:

(Price of milk balls ×quantity of milk balls) + (price of composite good ×Quantity of composite good)=Income

"2.50M+1C=150"

where M refers to quantity purchased of milk balls and C refers to quantity purchased of composite good. 

Now, we will find the intercept of both the axis using the below formula. 

Intercept of X-axis÷maximum quantity consumed of milk balls=income÷price of one milk ball

"=\frac{150}{2.50}\\=60"

Intercept of y-axis÷maximum quantity consumed of composite good=income÷price of one unit of composite good

"=\frac{150}{1}\\=150"

Now, we will take quantity of milk balls on x-axis and quantity of composite goods on the y-axis to plot the budget constraint. We have represented the budget constraint by line AB